Given f(x) = 5x2 – 3x + 1, find f(–2).
A) 15 B) 27 C) –13 D) –25

Answer by Nate(3500)

Given f(x) = 5x^2 – 3x + 1, find f(–2).
f(-2) = 5(-2)^2 – 3(-2) + 1
f(-2) = 5(4) + 6 + 1
f(-2) = 20 + 6 + 1
f(-2) = 27
(B)